# Galled Perfect Transfer Networks

### v1

@@ 18                   Add $(p_{N}(c_{min}),p_{N}(v))$ to $N$
21                   if  $c_{min}$ is marked as “visited” then
@@ 25                         Mark $v$ as “visited”
Let us now argue the complexity. Let $T=(V,E)$ be the given tree. The complexity will be dominated by the first for loop. For a given character $C\in\mathcal{C}$, computing the FAs can be done in time $O(|V|)$. Thus the first for can be done in time $O(|\mathcal{C}||V|)$. On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done on time $O(|V|+|\mathcal{C}|)$ since we know that every node can have at most $|\mathcal{C}|-1$ FA neighbors, as characters determine FA neighbors. Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g. [gusfield2014recombinatorics]). ∎
@@ 4 The Galled Compatibility Problem
Let $T$ be a tree and $v\in V(T)$. Let $C\in\mathcal{C}$ be a character and suppose that $C$ has two FAs $x_{1},x_{2}$ in $T$. In this case, we say that $C$ is split into $L_{T}(x_{1})$ and $L_{T}(x_{2})$ (noting that the union of these two leafsets must be $C$ since there are only two FAs). A character $C\in\mathcal{C}$ is maximal if there is no $C^{\prime}\in\mathcal{C}$ such that $C\subset C^{\prime}$. Two characters $A,B\in\mathcal{C}$ are compatible if there exists a tree $T$ in which $A$ and $B$ are clades, and incompatible otherwise. It is well-known that $A,B$ are compatible if and only if one of $A\cap B=\emptyset$ or one of $A$ or $B$ is a subset of the other. This means that if $A,B$ are incompatible, then the sets $A\cap B$, $A\setminus B,B\setminus A$ are all non-empty. Recall that for any set $\mathcal{C}$ of pairwise-compatible characters, there is a tree $T$ whose set of clades is exactly $\mathcal{C}$, plus the clade of the root and the leaves (see e.g. [gusfield2014recombinatorics]).
@@
The first direction of the if and only if statement is trivial: if $\mathcal{C}$ is galled-compatible, then any of its subset is galled-compatible, including $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, since a network that explains $\mathcal{C}$ also explains $\mathcal{C}_{1},\mathcal{C}_{2}$. In the other direction, suppose that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are galled-compatible. Since $C$ is compatible with every character, for any $A\in\mathcal{C}\setminus\{C\}$, one of $A\subseteq C,C\subseteq A$, or $A\cap C\neq\emptyset$ holds. Also, because $C$ is maximal, $C\subseteq A$ does not hold. It follows that $A$ is in one of $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ and thus $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ is a partition of $\mathcal{C}\setminus\{C\}$. Moreover, by defining $S_{1}:=\bigcup_{A\in\mathcal{C}_{1}}A$ and $S_{2}:=\bigcup_{B\in\mathcal{C}_{2}}B$, we get that $S_{1}\cap C=\emptyset$ and $S_{2}\subseteq C$. In particular, $S_{1}$ and $S_{2}$ are disjoint. Let $N_{1}$ and $N_{2}$ be galled trees that explain $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Then $L(N_{1})$ and $L(N_{2})$ are also disjoint. Consider the network $N$ obtained by (1) creating a root node $r$; (2) adding $r(N_{1})$ and $r(N_{2})$ as children of $r$; (3) for each $s\in C\setminus S_{2}$, adding $s$ as a leaf child of $r(N_{2})$. Observe that, since $N_{1}$ and $N_{2}$ are galled trees, by adding $r$ and its two child edges cannot create intersecting cycles, nor can adding leaves under $r(N_{2})$. Moreover, it is not hard to see that every character of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ is still explained by $N$, and $C$ is also explained since it is a clade of $N$. Thus $\mathcal{C}$ is galled-compatible. This construction also shows how to obtain $N$ from $N_{1}$ and $N_{2}$ in time $O(|C|)$, since it only requires adding a new root and adding up to $|C|$ leaves under $r(N_{2})$. ∎∎
@@
By incompatibility, $A$ and $B$ cannot both be clades of $T$. If only one of them is, say $A$ is a clade, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere. So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each. Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec b_{1}$ or $b_{1}\prec a_{1}$, and also $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$). The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles. ∎
@@
We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$. Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$. Since $b_{2}$ is incomparable with $b_{1}$, it is also incomparable with $a_{1}$. Then by Lemma 3 (part 1), $a_{2}=b_{2}$. Those imply $B_{1}\subseteq A_{1}$ and $A_{2}=B_{2}$, in turn implying $B\subseteq A$ and contradicting that $A$ and $B$ are incompatible.
@@
We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain. In Figure 5.3, $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$. One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$.
@@
Let $a_{1},a_{2}$ be the FAs of $A$, which respectively correspond to the clades $A_{1}$ and $A_{2}$. Let $X\in\mathcal{X}\setminus\{A\}$ and note that although $X$ intersects with both $A_{1}$ and $A_{2}$, it cannot contain both, as otherwise $A$ would not be maximal. Thus one of $X\cap A_{1}\subset A_{1}$ or $X\cap A_{2}\subset A_{2}$ holds.
@@
For an example, see the characters $B$ and $C$ in Figure 5. So far, we have handled characters “forced” by a split of $A$ into $A_{1}$ and $A_{2}$. As it turns out, the other characters can be handled in a recursive manner. To put this precisely, let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets. Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain. We say that a character $C\in\mathcal{C}$ is forced by $\{A_{1},A_{2}\}$ if either: $C\in\mathcal{X}$; $C$ contains the bottom of the $\mathcal{X}$ chain; or $C$ contains the stable side of the $\mathcal{X}$ chain. Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain.
@@
By Lemma 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$. Thus $T^{*}$ contains all the clades of $T$, plus possibly others. We assume that a clade present in both $T$ and $T^{*}$ is rooted at the same node in both trees — that is, if $y$ is the root of some clade $Y$ of $T$, then $y$ is also present in $T^{*}$ and is the root of clade $Y$ in $T^{*}$ as well. We will therefore assume that $V(T)\subseteq V(T^{*})$. Notably, $T^{*}$ also contains $X_{b}$ and $X_{s}$, so we assume that the same nodes $x_{b}$ and $x_{s}$ are the roots of these clades in $T^{*}$ as well.
@@
Now assume that a non-root node $v$ has both $x_{b}$ and $x_{s}$ as descendants. Let $a_{1}$ be the root of the $A_{1}$ clade. Note that $a_{1}$ is an ancestor of $x_{b}$ but not $x_{s}$, so $a_{1}$ is on the path from $x_{b}$ to $x_{s}$. This implies that $v$ is an ancestor of $a_{1}$ and thus the clade corresponding to $v$ contains $A_{1}$. As $v$ is also an ancestor of $x_{s}$, which roots the $A_{2}$ clade, then the clade of $v$ also contains $A_{2}$. Thus, the clade of $v$ contains $A_{1}\cup A_{2}=A$. By assumption, $A$ is not a clade of $T$, so the clade of $v$ is a strict superset of $A$. Moreover, since $v$ is not a root, the clade of $v$ must be in $T$ because it was forced, implying the existence of a character in $\mathcal{C}$ that strictly contains $A$, contradicting its maximality, thereby establishing our fact. ∎
@@
[…]
@@
Let $\mathcal{C}_{F}$ be the characters forced by $\{A_{1},A_{2}\}$. We will add further edges to $N$, but for the moment we claim that $N$, with the single transfer edge $(p_{N}(x_{b}),p_{N}(x_{s}))$, explains the characters in $\mathcal{C}_{F}$. Recall that there are three types of forced clades, which we deal with as follows:
@@
This summarizes the elements required for our algorithm. Due to space constraints, the full pseudo-code and analysis is relegated to the Appendix. However, its main steps can be summarized as follows:
  1. 1.
    If there is a maximal character compatible with every other, proceed recursively using Lemma 4.
  2. 2.
    Otherwise, find maximal incompatible characters $A,B$, guaranteed by Lemma 4.
    Initiate a tree whose clades are $A\cap B$ and $A\setminus B$.
  3. 3.
    Check whether Conditions 1-2-3 of Lemma 4 hold with respect to that partition. If so, consider the tree $T$ that contains all forced clades. For each $v\in V(T)$, list the set of characters $\mathcal{C}_{v}$ that only contain taxa whose parent is $v$. Solve for every such $\mathcal{C}_{v}$ recursively. If they all succeed, then we return that $\mathcal{C}$ is galled-compatible. If one of them fails, we may return that no solution is possible (because if one $\mathcal{C}_{v}$ is not galled-compatible, then $\mathcal{C}$ is not either).
  4. 4.
    If one of Condition 1-2-3 of Lemma 4 does not hold, then repeat the previous step but with a tree that contains clades $B\cap A,B\setminus A$. If this also fails, $\mathcal{C}$ is not galled-compatible.
Importantly, note that in this algorithm, we do not enter a recursion into $\mathcal{C}_{v}$ subsets for both partition attempts $\{A\cap B,A\setminus B\}$ and $\{B\cap A,B\setminus A\}$, as otherwise the complexity could be exponential. Since recursions are performed in at most one attempt of applying Step 3, we can argue that this takes polynomial time.
{toappendix}
Algorithm 2 is the full pseudocode of the galled-compatibility algorithm.
@@ 23       Let $T$ be the tree whose set of non-trivial clades is $F$
If no maximal compatible $C$ exists, the algorithm makes recursive calls on $\mathcal{C}_{v}$ sets in $tryPartition$, either when it receives $\{A_{1},A_{2}\}$ or $\{B_{1},B_{2}\}$. Suppose that recursive calls are made when $\{A_{1},A_{2}\}$ is received. Again, we observe that we either return “yes” or “no”, and then $getGalledTree$ exits without attempting $\{B_{1},B_{2}\}$. It thus suffices to count the recursive calls in one call of $tryPartition$, on disjoint subsets $\mathcal{C}_{v}$ of $\mathcal{C}$ (which are non-empty since this is checked by the algorithm). Note that these subsets do not contain $A$. Hence, the number of nodes in $R$ is at most

### v2

@@ 18                   Add $(p_{N}(c_{min}),p_{N}(v))$ to $N$
21                   if  $c_{min}$ is marked as “possibly simple” then
@@ 25                         Mark $v$ as “possibly simple”
Let us now argue the complexity. Let $T=(V,E)$ be the given tree. The complexity will be dominated by the first for loop. For a given character $C\in\mathcal{C}$, computing the FAs can be done in time $O(|V|)$. Thus the first for can be done in time $O(|\mathcal{C}||V|)$. On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done in time $O(|V|+|\mathcal{C}|)$. This is because we traverse $O(|V|)$ nodes, and each $\{u,v\}$ relationship of FA neighborhood requires work at most twice throughout the algorithm, once for $u$ and once for $v$. Moreover, each character implies at most one FA relationship, the total work over the FA neighbor hood relationships is $O(|\mathcal{C}|)$. We note that checking for the incomparability or descendance of nodes can be achieved in constant time with standard pre-processing of the tree. Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g. [gusfield2014recombinatorics]). ∎∎
@@ 4 The Galled Compatibility Problem
Let $T$ be a tree and $v\in V(T)$. Let $C\in\mathcal{C}$ be a character and suppose that $C$ has two FAs $x_{1},x_{2}$ in $T$. In this case, we say that $C$ is split into $L_{T}(x_{1})$ and $L_{T}(x_{2})$ (noting that the union of these two leafsets must be $C$ since there are only two FAs). A character $C\in\mathcal{C}$ is maximal if there is no $C^{\prime}\in\mathcal{C}$ such that $C\subset C^{\prime}$. Two characters $A,B\in\mathcal{C}$ are compatible if there exists a tree $T$ in which $A$ and $B$ are clades, and incompatible otherwise. It is well-known that $A,B$ are compatible if and only if either $A\cap B=\emptyset$ or one of $A$ or $B$ is a subset of the other. This means that if $A,B$ are incompatible, then the sets $A\cap B$, $A\setminus B,B\setminus A$ are all non-empty. Recall that for any set $\mathcal{C}$ of pairwise-compatible characters, there is a tree $T$ whose set of clades is exactly $\mathcal{C}$, plus the clade of the root and the leaves (see e.g. [gusfield2014recombinatorics]).
@@
The first direction of the if and only if statement is trivial: if $\mathcal{C}$ is galled-compatible, then any of its subset is galled-compatible, including $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, since a network that explains $\mathcal{C}$ also explains $\mathcal{C}_{1},\mathcal{C}_{2}$. In the other direction, suppose that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are galled-compatible. Since $C$ is compatible with every character, for any $A\in\mathcal{C}\setminus\{C\}$, one of $A\subseteq C,C\subseteq A$, or $A\cap C\neq\emptyset$ holds. Also, because $C$ is maximal, $C\subseteq A$ does not hold. It follows that $A$ is in one of $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ and thus $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ is a partition of $\mathcal{C}\setminus\{C\}$. Moreover, by defining $S_{1}:=\bigcup_{A\in\mathcal{C}_{1}}A$ and $S_{2}:=\bigcup_{B\in\mathcal{C}_{2}}B$, we get that $S_{1}\subseteq C$ and $S_{2}\cap C=\emptyset$. In particular, $S_{1}$ and $S_{2}$ are disjoint. Let $N_{1}$ and $N_{2}$ be galled trees that explain $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Then $L(N_{1})$ and $L(N_{2})$ are also disjoint. Consider the network $N$ obtained by (1) creating a root node $r$; (2) adding $r(N_{1})$ and $r(N_{2})$ as children of $r$; (3) for each $s\in C\setminus S_{1}$, adding $s$ as a leaf child of $r(N_{2})$. Observe that, since $N_{1}$ and $N_{2}$ are galled trees, by adding $r$ and its two child edges cannot create intersecting cycles, nor can adding leaves under $r(N_{2})$. Moreover, it is not hard to see that every character of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ is still explained by $N$, and $C$ is also explained since it is a clade of $N$. Thus $\mathcal{C}$ is galled-compatible. This construction also shows how to obtain $N$ from $N_{1}$ and $N_{2}$ in time $O(|C|)$, since it only requires adding a new root and adding up to $|C|$ leaves under $r(N_{2})$. ∎∎
@@
By incompatibility, $A$ and $B$ cannot both be clades of $T$. If only one of them is a clade, say $A$, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere. So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each. Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec b_{1}$ or $b_{1}\prec a_{1}$, and also $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$). The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles. ∎
@@
We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$. Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$. Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{1}$ is also incomparable with $a_{1}$. Then by Lemma 3 (part 1), $a_{2}=b_{2}$. Those imply $B_{1}\subseteq A_{1}$ and $A_{2}=B_{2}$, in turn implying $B\subseteq A$ and contradicting that $A$ and $B$ are incompatible.
@@
We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain. In Figure 5.3, $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$. One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$. The concept of a chain is illustrated in Figure 7.
Figure 7: An $(A_{1},A_{2})$-chain, where $X_{1},X_{2},X_{3}\in\mathcal{X}$ all intersect with both $A_{1}$ and $A_{2}$. On the $A_{1}$ side, the FAs are ordered by ancestry, and the only way to explain all these characters is to have them use the same transfer to reach the $A_{2}$ side.
@@
Let $a_{1},a_{2}$ be the FAs of $A$ in $T$, which respectively correspond to the clades $A_{1}$ and $A_{2}$. Let $X\in\mathcal{X}\setminus\{A\}$ and note that although $X$ intersects with both $A_{1}$ and $A_{2}$, it cannot contain both, as otherwise $A$ would not be maximal. Thus one of $X\cap A_{1}\subset A_{1}$ or $X\cap A_{2}\subset A_{2}$ holds.
@@
For an example, see the characters $B$ and $C$ in Figure 5. So far, we have handled characters “forced” by a split of $A$ into $A_{1}$ and $A_{2}$. As it turns out, the other characters can be handled in a recursive manner.
To put this precisely, let us gather all the information on the tree that has been shown to be forced so far. Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets. Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain. We say that a character $C\in\mathcal{C}$ is forced by $\{A_{1},A_{2}\}$ if either: $C\in\mathcal{X}$; $C$ contains the bottom of the $\mathcal{X}$ chain; or $C$ contains the stable side of the $\mathcal{X}$ chain.
Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain. The first type of clades are enforced due to Lemma 4, and the second type due to Lemma 4.
@@
By Lemma 4 and Lemma 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$. Thus $T^{*}$ contains all the clades of $T$, plus possibly others. We assume that a clade present in both $T$ and $T^{*}$ is rooted at the same node in both trees — that is, if $y$ is the root of some clade $Y$ of $T$, then $y$ is also present in $T^{*}$ and is the root of clade $Y$ in $T^{*}$ as well. We will therefore assume that $V(T)\subseteq V(T^{*})$. Notably, $T^{*}$ also contains $X_{b}$ and $X_{s}$, so we assume that the same nodes $x_{b}$ and $x_{s}$ are the roots of these clades in $T^{*}$ as well.
@@
Now assume that a non-root node $v$ has both $x_{b}$ and $x_{s}$ as descendants in $T$. Let $a_{1}$ be the root of the $A_{1}$ clade. Note that $a_{1}$ is an ancestor of $x_{b}$ but not $x_{s}$, so $a_{1}$ is on the path from $x_{b}$ to $x_{s}$ (when viewing $T$ as an undirected graph). This implies that $v$ is an ancestor of $a_{1}$ and thus the clade corresponding to $v$ contains $A_{1}$. As $v$ is also an ancestor of $x_{s}$, which roots the $A_{2}$ clade, then the clade of $v$ also contains $A_{2}$. Thus, the clade of $v$ contains $A_{1}\cup A_{2}=A$. By assumption, $A$ is not a clade of $T$, so the clade of $v$ is a strict superset of $A$. Moreover, since $v$ is not a root, the clade of $v$ must be in $T$ because it was forced, implying the existence of a character in $\mathcal{C}$ that strictly contains $A$, contradicting its maximality, thereby establishing our fact. ∎
@@
Proof Since $X_{1}$ is split into $X_{b}$ and $X_{s}$ in $T^{*}$, by Lemma 3 there must be a transfer edge between $p_{N}(x_{b})$ or a descendant of $x_{b}$, and $p_{N}(x_{s})$ and a descendant of $x_{s}$, which implies that existence of the claimed cycle. ∎
@@
Let $\mathcal{C}_{F}$ be the characters forced by $\{A_{1},A_{2}\}$. We will add further edges to $N$, but for the moment we claim that $N$, with the single transfer edge $(p_{N}(x_{b}),p_{N}(x_{s}))$, explains the characters in $\mathcal{C}_{F}$. Recall that there are three types of forced characters, which we deal with as follows:
@@
Algorithm 2 is the full pseudocode of the galled-compatibility algorithm. For simplicity, it only solves the decision version of the problem (i.e., whether the characters are galled-compatible or not), but it can easily be adapted to reconstruct a network.
@@ 23       Let $T$ be the tree whose set of non-trivial clades is $F$
If no maximal compatible $C$ exists, the algorithm makes recursive calls on $\mathcal{C}_{v}$ sets in $tryPartition$, either when it receives $\{A_{1},A_{2}\}$ or $\{B_{1},B_{2}\}$ (but not both). Suppose that recursive calls are made when $\{A_{1},A_{2}\}$ is received. Again, we observe that we either return “yes” or “no”, and then $getGalledTree$ exits without attempting $\{B_{1},B_{2}\}$. It thus suffices to count the recursive calls in one call of $tryPartition$, on disjoint subsets $\mathcal{C}_{v}$ of $\mathcal{C}$ (which are non-empty since this is checked by the algorithm). Note that these subsets do not contain $A$. Hence, the number of nodes in $R$ is at most